# A priori error analysis of the proximal Galerkin method

### v2

@@ 4.1. Preliminaries for the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$-element pair
The subspaces given in 94 satisfies the inf-sup condition 40 and there exists a stable Fortin operator satisfying 42 and
@@
The proof is based on the arguments in [36, Appendix B]. The operator $\Pi_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is constructed as follows:
@@
where $\tilde{\mathcal{I}}_{h}:L^{1}(\Omega)\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{0}(\Omega)$ is the quasi-interpolant introduced in [23, Section 6], $I$ is the identity operator, and $\tilde{\Pi}_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is defined element-wise to satisfy $(\tilde{\Pi}_{h}v)|_{T}:=b_{T}v_{T}\in\mathbb{B}(T)$, where $v_{T}\in\mathbb{P}_{0}(T)=\mathbb{R}$ solves
@@
It is easy to see that $v_{T}=(v,1)_{T}/(b_{T},1)_{T}$ is the unique solution to the above. Owing to mesh regularity and to the properties of the bubble function [49, Lemma 4.1], we have
@@
Using 102 in 99 along with the triangle inequality and the stability of $\tilde{\mathcal{I}}_{h}$, we obtain the stability of $\Pi_{h}$ in $L^{2}(\Omega)$. To show stability in the $H^{1}$-semi norm, we apply triangle inequality, a local inverse estimate, 102 and the stability and approximation properties of $\tilde{\mathcal{I}}_{h}$ [23, Lemma 6.3 and Theorem 6.4]:
@@
(104) $K_{h}\subset\left\{v_{h}\in V_{h}\mid\int_{T}(v_{h}-\phi)\operatorname{d}\!x\geq 0\;~\text{ for all }T\in\mathcal{T}_{h}\right\}.$
To satisfy Assumption 2, we first require a map defined over $V$ and mapping the set $K_{h}$ to the set $K$. We will later shift this map to construct the enriching map $\mathcal{E}_{h}$, see 121. The natural choice is the Clément interpolant [17, 5]. Here, we utilize a specialized variant of this interpolant that employs weighted local averages over vertex patches to maintain second-order accuracy for smooth functions. This specific interpolant was first introduced in [29], but we slightly modify its definition to incorporate non-homogeneous boundary data.
@@
The following is a modified proof of [29, Theorem 11] for the case of inhomogeneous boundary data and $H^{1+s}(\Omega)$ functions. The key observation, as we show next, is that this interpolant preserves linear functions locally [29]; i.e., $\mathcal{C}_{h}p=p$ for any $p\in\mathbb{P}_{1}(\omega_{z})$ and $z\in\mathcal{N}_{h}$. For an interior node $z\in\mathcal{N}_{h}\backslash\partial\Omega$, it is clear that $b_{z}=b$ for a constant $b$ since $\sum_{T\subset\omega_{z}}\alpha_{z,T}=1$ and for a linear polynomial $p$, it follows that:
@@
In the above, we also used that $\mathrm{card}(\omega_{z})$ is uniformly bounded w.r.t. $h$ for all $z\in\mathcal{N}_{h}$, which follows from the shape-regularity of the mesh; see, e.g., [24, Proposition 11.6]. If $z\in\partial\Omega$, then $z\in\partial T\cap\omega_{z}$ for some $T\subset\omega_{z}$. Using the definition of $\mathcal{C}_{h}$ on boundary nodes, we apply local inverse and trace inequalities to bound
@@
In the last line above, we used the local stability property of $\mathcal{SZ}_{h}$ [46, Theorem 3.1]. Using that $C_{h}(p-w)|_{T}=\sum_{z\in\mathcal{V}_{T}}C_{h}(p-w)(z)\phi_{z}$ and that $\|\phi_{z}\|_{L^{2}(T)}\lesssim h_{T}^{n/2}$, we arrive at
@@
Combining 109 with 112 and using the Bramble–Hilbert Lemma (see, e.g., [23, Lemma 5.6]) yields the required bound on the first term of 107. To obtain the required bound on the second term, we apply triangle and inverse estimates:
@@
The proof is completed by using the approximation properties of $\mathcal{SZ}_{h}$ [46, Theorem 4.1] and the proven bound on $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}$. ∎
@@ 4.2. Preliminaries for the $(\mathbb{P}_{1},\mathbb{P}_{1})$-element pair
For this case, the Fortin operator is simply the $L^{2}(\Omega)$-projection; cf. [36, Remark 5.2]. Thus, the main task is to construct the enriching map of Assumption 2, which will again depend on a modified Clément quasi-interpolant. Using the definition of $K_{h}$ 74 and testing with $\varphi_{z}$, we observe that
(113) $K_{h}\subset\left\{v_{h}\in V_{h}\mid\int_{\omega_{z}}(v_{h}-\phi)\varphi_{z}\operatorname{d}\!x\geq 0~\text{ for all }z\in\mathcal{N}_{h}\backslash\partial\Omega\right\},$
@@
where we used that $\int_{T}\varphi_{z^{\prime}}\varphi_{z}\operatorname{d}\!x=(n+1)^{-1}(n+2)^{-1}|T|(1+\delta_{z^{\prime},z})$ [15, Exercise 4.1.1]. Hence,
@@ 4.3. Error estimates for the obstacle problem
The preliminary results in Sections 4.1 and 4.2 leave us ready to state and prove optimal a priori error estimates for the PG method applied to the obstacle problem.
@@
We verify the assumptions of Theorem 3.6. For the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ elements, Lemma 4.1 shows that 76a-76b and 78a in Assumption 3 hold. For the $(\mathbb{P}_{1},\mathbb{P}_{1})$ elements, the Fortin operator is simply the $L^{2}(\Omega)$ projection onto $V_{h}$. It is standard to show that 76a-76b and 78a hold if the mesh is quasi-uniform, as assumed in this case; see, e.g., [24, Proposition 22.21] and [36, Remark 5.2].
To verify Assumption 2, define the enriching map as follows:
@@
To show 120, we utilize Lemma 3.5 and note that $P_{W_{h}}\tilde{u}_{h}^{\ell}=\tilde{u}_{h}^{\ell}$ since $\tilde{u}_{h}^{\ell}\in\mathbb{P}_{0}(\mathcal{T}_{h})$. Therefore, since $B$ is the identity map and $\Omega_{d}=\Omega$ in this case, 71 reads
(123) $\|u^{*}-\tilde{u}_{h}^{\ell}\|_{L^{2}(\Omega)}$ $\leq\|u_{h}^{\ell}-u^{*}\|_{L^{2}(\Omega)}+\|(I-P_{W_{h}})u_{h}^{\ell}\|_{L^{2}(\Omega)}$
$\leq\|u_{h}^{\ell}-u^{*}\|_{L^{2}(\Omega)}+\|(I-P_{W_{h}})(u_{h}^{\ell}-u^{*})\|_{L^{2}(\Omega)}+\|(I-P_{W_{h}})u^{*}\|_{L^{2}(\Omega)}$
@@
where we used the stability of $P_{W_{h}}$ and that $\|u^{*}-P_{W_{h}}u^{*}\|_{L^{2}(T)}\lesssim h_{K}|u^{*}|_{H^{1}(T)}$ for any $T\in\mathcal{T}_{h}$. The final estimate 120 is obtained by combining 119 and the final inequality in 123. ∎
@@ 5. Application II: The Signorini problem
We provide Algorithm 3, the application of the PG method (Algorithm 1) to this problem for completeness; see also [19, Example 2].
@@
Main goal: We derive error estimates for Algorithm 3 in Corollary 3. We also apply the framework presented in Section 3, utilizing Theorem 3.6. To this end, we proceed by constructing Fortin and enriching operators satisfying Assumptions 2 and 3.
@@
In addition, the Fortin operator $\Pi_{h}:H^{1}_{\mathrm{D}}(\Omega)^{2}\rightarrow V_{h}$ given by $\Pi_{h}w=(\tilde{\Pi}_{h}w_{n})n+(\mathcal{SZ}_{h}w_{\tau})\tau$ satisfies 42 and the stability and approximation bounds 76a, 76b and 78a stated in Assumption 3.
@@
where we used the stability of $\mathcal{SZ}_{h}$ in the $H^{1}(\Omega)$-seminorm, see [8, Corollary 4.8.15]. The final error estimate for $\tilde{\Pi}_{h}$ in the $H^{1}(\Omega)$-seminorm is proven similarly. ∎
In order to check Assumption 2, we observe that the set $K_{h}$ 74 satisfies
(136) $K_{h}\subset\{u_{h}\in V_{h}\mid(u_{h}\cdot n-g,\varphi_{z})_{\Gamma_{\mathrm{T}}}\leq 0~\text{ for all }z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}\}.$
To define the enriching map required by Assumption 2, we begin by defining a modified Clément interpolant ${\mathcal{C}}_{h}:H^{1}(\Omega)\rightarrow H^{1}(\Omega)\cap\mathbb{P}_{1}(\mathcal{T}_{h})$ using the strategy outlined in Section 4.2. To this end, we require a technical assumption on the mesh. Note that this assumption can be easily satisfied by a local refinement near the boundary $\partial\Gamma_{\mathrm{T}}=\overline{\Gamma_{\rm T}}\cap\overline{\Gamma_{\rm D}}$.
@@
We now define ${\alpha}_{z,i}$. Let $(\hat{w}_{j},\hat{x}_{j})$, $j=0,1$, be the non-negative weights and points given by the Gauss–Radau quadrature rule [24, Table 6.1], which is an exact quadrature rule for polynomials of degree $2$ on the reference facet $\hat{E}$. Let $F_{E}$ denote the affine linear transformation from the reference facet $\hat{E}$ to a facet $E\subset\Gamma_{\mathrm{T}}$. Observe that these quadrature points define weights and points $(w_{j},x_{j})$ with $x_{j}\in\tilde{\omega}_{z}$ for $j=0,1,2,3$, such that for any continuous piecewise second-order polynomial $v$ on $\tilde{\omega}_{z}=E_{1}\cup E_{2}$, we can write
@@ Acknowledgments
We thank Ioannis P. A. Papadopoulos for his comments on the manuscript and Thomas M. Surowiec and Dohyun Kim for numerous interesting discussions. BK and RM were supported in part by the U.S. Department of Energy Office of Science, Early Career Research Program under Award Number DE-SC0024335 and by the Center for Information Geometric Mechanics (CIGMO), a PSAAP-IV Focused Investigatory Center funded by the Department of Energy under Award Number DE-NA0004261. Keith’s work was also supported in part by the Alfred P. Sloan Foundation via a Sloan Research Fellowship in Mathematics. MZ acknowledges support from an ETH Postdoctoral Fellowship for the project “Reliable, Efficient, and Scalable Methods for Scientific Machine Learning”.

### v3

@@ 4.1. Preliminaries for the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$-element pair
The subspaces given in 94 satisfies the inf-sup condition 41 and there exists a stable Fortin operator satisfying 43 and
@@
The proof is based on the arguments in [40, Appendix B]. The operator $\Pi_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is constructed as follows:
@@
where $\tilde{\mathcal{I}}_{h}:L^{1}(\Omega)\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{0}(\Omega)$ is the quasi-interpolant introduced in [25, Section 6], $I$ is the identity operator, and $\tilde{\Pi}_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is defined element-wise to satisfy $(\tilde{\Pi}_{h}v)|_{T}:=b_{T}v_{T}\in\mathbb{B}(T)$, where $v_{T}\in\mathbb{P}_{0}(T)=\mathbb{R}$ solves
@@
It is easy to see that $v_{T}=(v,1)_{T}/(b_{T},1)_{T}$ is the unique solution to the above. Owing to mesh regularity and to the properties of the bubble function [53, Lemma 4.1], we have
@@
Using 102 in 99 along with the triangle inequality and the stability of $\tilde{\mathcal{I}}_{h}$, we obtain the stability of $\Pi_{h}$ in $L^{2}(\Omega)$. To show stability in the $H^{1}$-semi norm, we apply triangle inequality, a local inverse estimate, 102 and the stability and approximation properties of $\tilde{\mathcal{I}}_{h}$ [25, Lemma 6.3 and Theorem 6.4]:
@@
(104) $K_{h}=\left\{v_{h}\in V_{h}\mid\int_{T}(v_{h}-\phi)\operatorname{d}\!x\geq 0\;~\text{ for all }T\in\mathcal{T}_{h}\right\}.$
To satisfy Assumption 2, we first require a map defined over $V$ and mapping the set $K_{h}$ to the set $K$. We will later shift this map to construct the reconstruction operator $\mathcal{E}_{h}$, see 121. The natural choice is the Clément interpolant [19, 6]. Here, we utilize a specialized variant of this interpolant that employs weighted local averages over vertex patches to maintain second-order accuracy for smooth functions. This specific interpolant was first introduced in [31], but we slightly modify its definition to incorporate non-homogeneous boundary data.
@@
The following is a modified proof of [31, Theorem 11] for the case of inhomogeneous boundary data and $H^{1+s}(\Omega)$ functions. The key observation, as we show next, is that this interpolant preserves linear functions locally [31]; i.e., $\mathcal{C}_{h}p=p$ for any $p\in\mathbb{P}_{1}(\omega_{z})$ and $z\in\mathcal{N}_{h}$. For an interior node $z\in\mathcal{N}_{h}\backslash\partial\Omega$, it is clear that $b_{z}=b$ for a constant $b$ since $\sum_{T\subset\omega_{z}}\alpha_{z,T}=1$ and for a linear polynomial $p$, it follows that:
@@
In the above, we also used that $\mathrm{card}(\omega_{z})$ is uniformly bounded w.r.t. $h$ for all $z\in\mathcal{N}_{h}$, which follows from the shape-regularity of the mesh; see, e.g., [26, Proposition 11.6]. If $z\in\partial\Omega$, then $z\in\partial T\cap\omega_{z}$ for some $T\subset\omega_{z}$. Using the definition of $\mathcal{C}_{h}$ on boundary nodes, we apply local inverse and trace inequalities to bound
@@
In the last line above, we used the local stability property of $\mathcal{SZ}_{h}$ [50, Theorem 3.1]. Using that $C_{h}(p-w)|_{T}=\sum_{z\in\mathcal{V}_{T}}C_{h}(p-w)(z)\phi_{z}$ and that $\|\phi_{z}\|_{L^{2}(T)}\lesssim h_{T}^{n/2}$, we arrive at
@@
Combining 109 with 112 and using the Bramble–Hilbert Lemma (see, e.g., [25, Lemma 5.6]) yields the required bound on the first term of 107. To obtain the required bound on the second term, we apply triangle and inverse estimates:
@@
The proof is completed by using the approximation properties of $\mathcal{SZ}_{h}$ [50, Theorem 4.1] and the proven bound on $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}$. ∎
@@ 4.2. Preliminaries for the $(\mathbb{P}_{1},\mathbb{P}_{1})$-element pair
For this case, the Fortin operator is simply the $L^{2}(\Omega)$-projection; cf. [40, Remark 5.2]. Thus, the main task is to construct the reconstruction operator associated with Assumption 2, which again depends on a modified Clément quasi-interpolant. Using the definition of $K_{h}$ 75b and testing with $\varphi_{z}$, we observe that
(113) $K_{h}=\left\{v_{h}\in V_{h}\mid\int_{\omega_{z}}(v_{h}-\phi)\varphi_{z}\operatorname{d}\!x\geq 0~\text{ for all }z\in\mathcal{N}_{h}\backslash\partial\Omega\right\},$
@@
where we used that $\int_{T}\varphi_{z^{\prime}}\varphi_{z}\operatorname{d}\!x=(n+1)^{-1}(n+2)^{-1}|T|(1+\delta_{z^{\prime},z})$ [17, Exercise 4.1.1]. Hence,
@@ 4.3. Error estimates for the obstacle problem
The preliminary results in Sections 4.1 and 4.2 leave us ready to state and prove optimal a priori error estimates for the PG method applied to the obstacle problem. We emphasize that the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ pair does not require quasi-uniform meshes.
@@
We verify the assumptions of Theorem 3.6. For the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ elements, Lemma 4.1 shows that 76a-76b and 78a in Assumption 3 hold. For the $(\mathbb{P}_{1},\mathbb{P}_{1})$ elements, the Fortin operator is simply the $L^{2}(\Omega)$ projection onto $V_{h}$. It is standard to show that 76a-76b and 78a hold if the mesh is quasi-uniform, as assumed in this case; see, e.g., [26, Proposition 22.21] and [40, Remark 5.2].
To verify Assumption 2, define the reconstruction operator as follows:
@@
To show 120, we utilize Lemma 3.5 and note that $P_{h}\tilde{u}_{h}^{\ell}=\tilde{u}_{h}^{\ell}$ since $\tilde{u}_{h}^{\ell}\in\mathbb{P}_{0}(\mathcal{T}_{h})$. Therefore, since $B$ is the identity map and $\Omega_{d}=\Omega$ in this case, 72 reads
(123) $\|u^{*}-\tilde{u}_{h}^{\ell}\|_{L^{2}(\Omega)}$ $\leq\|u_{h}^{\ell}-u^{*}\|_{L^{2}(\Omega)}+\|(I-P_{h})u_{h}^{\ell}\|_{L^{2}(\Omega)}$
$\leq\|u_{h}^{\ell}-u^{*}\|_{L^{2}(\Omega)}+\|(I-P_{h})(u_{h}^{\ell}-u^{*})\|_{L^{2}(\Omega)}+\|(I-P_{h})u^{*}\|_{L^{2}(\Omega)}$
@@
where we used the stability of $P_{h}$ and that $\|u^{*}-P_{h}u^{*}\|_{L^{2}(T)}\lesssim h_{K}|u^{*}|_{H^{1}(T)}$ for any $T\in\mathcal{T}_{h}$. The final estimate 120 is obtained by combining 119 and the final inequality in 123. ∎
@@ 5. Application II: The Signorini problem
We provide Algorithm 3, the application of the PG method (Algorithm 1) to this problem for completeness; see also [21, Example 2].
@@
Main goal: We derive error estimates for Algorithm 3 in Corollary 3. We also apply the framework presented in Section 3, utilizing Theorem 3.6. To this end, we proceed by constructing Fortin and reconstruction operators satisfying Assumptions 2 and 3.
@@
In addition, the Fortin operator $\Pi_{h}:H^{1}_{\mathrm{D}}(\Omega)^{2}\rightarrow V_{h}$ given by $\Pi_{h}w=(\tilde{\Pi}_{h}w_{n})n+(\mathcal{SZ}_{h}w_{\tau})\tau$ satisfies 43 and the stability and approximation bounds 76a, 76b and 78a stated in Assumption 3.
@@
where we used the stability of $\mathcal{SZ}_{h}$ in the $H^{1}(\Omega)$-seminorm, see [9, Corollary 4.8.15]. The final error estimate for $\tilde{\Pi}_{h}$ in the $H^{1}(\Omega)$-seminorm is proven similarly. ∎
In order to check Assumption 2, we observe that the set $K_{h}$ 75b satisfies
(136) $K_{h}=\{u_{h}\in V_{h}\mid(u_{h}\cdot n-g,\varphi_{z})_{\Gamma_{\mathrm{T}}}\leq 0~\text{ for all }z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}\}.$
To define the reconstruction operator required by Assumption 2, we begin by defining a modified Clément interpolant ${\mathcal{C}}_{h}:H^{1}(\Omega)\rightarrow H^{1}(\Omega)\cap\mathbb{P}_{1}(\mathcal{T}_{h})$ using the strategy outlined in Section 4.2. To this end, we require a technical assumption on the mesh. Note that this assumption can be easily satisfied by a local refinement near the boundary $\partial\Gamma_{\mathrm{T}}=\overline{\Gamma_{\rm T}}\cap\overline{\Gamma_{\rm D}}$.
@@
We now define ${\alpha}_{z,i}$. Let $(\hat{w}_{j},\hat{x}_{j})$, $j=0,1$, be the non-negative weights and points given by the Gauss–Radau quadrature rule [26, Table 6.1], which is an exact quadrature rule for polynomials of degree $2$ on the reference facet $\hat{E}$. Let $F_{E}$ denote the affine linear transformation from the reference facet $\hat{E}$ to a facet $E\subset\Gamma_{\mathrm{T}}$. Observe that these quadrature points define weights and points $(w_{j},x_{j})$ with $x_{j}\in\tilde{\omega}_{z}$ for $j=0,1,2,3$, such that for any continuous piecewise second-order polynomial $v$ on $\tilde{\omega}_{z}=E_{1}\cup E_{2}$, we can write
@@ Acknowledgments
We thank Ioannis P. A. Papadopoulos for his comments on the manuscript and Thomas M. Surowiec and Dohyun Kim for numerous interesting discussions. BK and RM were supported in part by the U.S. Department of Energy, Office of Science Early Career Research Program under Award Number DE-SC0024335 and by the Center for Information Geometric Mechanics and Optimization (CIGMO), a PSAAP-IV Focused Investigatory Center funded by the U.S. Department of Energy, National Nuclear Security Administration under Award Number DE-NA0004261. Keith’s work was also supported in part by the Alfred P. Sloan Foundation via a Sloan Research Fellowship in Mathematics. MZ acknowledges support from an ETH Postdoctoral Fellowship for the project “Reliable, Efficient, and Scalable Methods for Scientific Machine Learning”.